OpenStudy (anonymous):

Write the slope-intercept form of the line which passes through (4,2) with a slope of 1/2. A.y=1/2 B.y=1/2x+4

OpenStudy (anonymous):

so slope = 1/2 and points are 4,2. using the values in y = mx+c, 2 = 1/2. 4 + c which gives c = 0. so equation of line = y = 1/2 x.